OpenStudy (anonymous):

from 1970 to 2011, housing prices increased by an annualized rate of 4.89%. The median price for a house in 1970 was about $27000. what is the median price for a house in 2011?

OpenStudy (anonymous):

ok so it is 41 years later, use \[27000\times (1.0489)^{41}\] with a calculator

OpenStudy (anonymous):

i get \[\$191,188.28\] rounded

OpenStudy (anonymous):

omg, so far you have gotten all the right answersssss. Ahhhh how are you sooo smart hah

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

how about 5^(x-3)=137

OpenStudy (anonymous):

\[x-3=\frac{\ln(135)}{\ln(5)}\] \[x=\frac{\ln(135)}{\ln(5)}+3\]

OpenStudy (anonymous):

hold on typo there

OpenStudy (anonymous):

\[x=\frac{\ln(137)}{\ln(5)}+3\]

OpenStudy (anonymous):

how do you know when to use ln, and when to use log

OpenStudy (anonymous):

3^(x/7)=0.2another one lol

OpenStudy (anonymous):

ok if i see \[b^x=A\] and i want x, i write \[x=\frac{\ln(A)}{\ln(b)}\] the log of the total divided by the log of the base.

OpenStudy (anonymous):

guess what? \[\frac{x}{7}=\frac{\ln(.2)}{\ln(3)}\] just like above

OpenStudy (anonymous):

how do you get rid of the 7 though

OpenStudy (anonymous):

i got -10.254. is that correct?